# PADE: A Predictor-Free Sparse Attention Accelerator via Unified Execution and Stage Fusion

### v1

@@ V-PU A Systolic Array with 8 $\times$ 16 INT8 PEs;
GPU comparison We benchmark on an Nvidia H100 using SOTA TensorRT-LLM [Nvidia2023] with FlashAttention3 [shah2024flashattention]. To exclude the software overhead, we measure execution time with cudaEvent, isolating GPU execution from CPU interference. The GPU is dedicated during testing, and large batch sizes are used to amortize data transfer costs. Non-computational phases are excluded using nvprof. Power is measured via nvidia-smi; dynamic power is computed as the difference between active and idle states. Each experiment is run 2k times, discarding the top and bottom $15\%$ before averaging.
@@ VI-B Algorithm Performance
Fig. 15 (a)(b) compares the accuracy of three representative software-only sparse attention methods, two predictor-free works (SpAtten [wang2021spatten], DTATrans [yang2022dtatrans]), and PADE. The Sparsity Level denotes the ratio between the computation cost of sparse execution (prediction + actual computation) and that of dense execution. StreamingLLM [xiao2023efficient] adopts a static sparsity pattern—retaining only the initial and recent tokens, while MInference [jiang2024minference], and DoubleSparsity [yang2024post] rely on runtime sparsity prediction. Four key observations are made: 1) StreamingLLM performs the worst, due to its lack of adaptivity in capturing input-dependent sparsity patterns. 2) MInference improves accuracy by combining dynamic prediction with predefined sparsity patterns but remains limited by restricted pattern diversity. 3) DoubleSparsity introduces a more flexible dynamic sparsity mechanism and integrates channel sparsity to reduce prediction overhead. However, as its prediction computation and memory access cannot be reused in subsequent steps, it suffers from inefficiency, yielding slightly lower accuracy than PADE at the same sparsity level. 4) For SpAtten and DTATrans, comparable accuracy to PADE can be achieved only after fine-tuning. This is mainly because both methods eliminate the predictor by using the previous layer’s attention distribution to guide pruning in the current layer, which introduces significant errors without fine-tuning. In contrast, PADE consistently achieves the best performance across all settings, benefiting from its fine-grained sparsity prediction and efficient reuse of both computation and memory access.
@@ VI-E Comparison to SOTA Accelerators
Workload Balance. Fig. 23 (a) presents a detailed breakdown of execution time across varying numbers of PE lanes to illustrate load balance effects. To better show PADE’s advantages, we compare it against a SOTA bit accelerator, BitWave [shi2024bitwave], which leverages bit-flipping to enhance bit-plane sparsity. Since each PE lane integrates multiple bit-serial multipliers, intra-PE stall arises when certain multipliers handle more effective bits. Inter-PE install, in contrast, results from variations in bit sparsity across different key vectors. As the number of PE lanes increase, BitWave suffers from greater intra- and inter-PE imbalance, as it only exploits bit-0 sparsity, which exhibits large variability. In comparison, PADE adopts a more balanced bit sparsity distribution, thereby achieving around $30\%$ higher PE utilization.
@@ VI-F Discussion for Deployment and Scalability
Extension for MXINT. The micro-scale format performs fine-grained quantization along the channel dimension by grouping data into 32-element segments [rouhani2023microscaling]. PADE ensures compatibility by applying group-wise scaling to the bit uncertain interval (BUI) (§IV-A). As shown in Fig. 25 (a), when processing 64-length Q and K vectors, the micro-format partitions them into two 32-element groups and quantizes each group using calibration-derived factors ($\Delta$), resulting in group-wise BUI scaling. As depicted in Fig. 25 (b), PADE achieves compatibility via two steps. ❶ First, it performs bitwise serial multiplication within each 32-length group, with the BUI derived from the strategy in §IV-A. The resulting BUI is then expanded using the quantization factor, eg., $\Delta_{\rm Q1}\Delta_{\rm K1}/\Delta_{\rm A}$. ❷ Finally, the max and min BUI values across all groups are aggregated to compute the overall BUI. This method can be extended to dot products of arbitrary length.
Extension for FP formats. Prior studies [sheng2023flexgen, xiao2023smoothquant, zhao2024atom] have demonstrated that the K and V tensors are highly amenable to quantization because the softmax normalization in self-attention naturally suppresses quantization noise. This property enables safe quantization to INT8 or even INT4 with negligible accuracy degradation. Motivated by this, when queries operate in FP format, PADE converts the INT-FP computation into a bit-serial form through exponent alignment, following methodologies adopted in prior works [chen2025bitmod, jang2024figna, fang2025anda]. The resulting bit-serial execution is fully compatible with PADE’s existing processing mechanism.
@@ VI-G Limitations and Future Direction
While PADE significantly advances sparse-attention acceleration by eliminating the sparsity predictor through unified bit-serial stage fusion, several limitations remain for future works: (1) As the sequence length of large language models (LLMs) continues to scale, the need for distributed attention becomes inevitable. Extending PADE to distributed scenarios, especially in emerging wafer-scale architectures [wang2026WATOS, wang2026TEMP, wang2024tmac, he2025waferllm, hu2024wafer, tang2025moentwine], remains an open problem. (2) Beyond single-bit granularity, exploring multi-bit stage fusion may provide a more favorable efficiency–accuracy trade-off, which represents a promising direction for future PADE enhancements.
@@ VII Related Works
Efficient Attention Accelerators. Numerous attention accelerators [wang2025bitstopper, wang2025lapa, wang2025beta, li2020ftrans, ham20203, ham2021elsa, qu2022dota, fan2023taskfusion, yang2022dtatrans, hong2022dfx, liu2024hsconn, wang2024sofa, wang2022energy, fang2022algorithm, yazdanbakhsh2022sparse, li2022accelerating, lu2021sanger, wang2021spatten, zhou2022energon, qin2023fact, shen2022salo, fan2022adaptable, bai2024swat, zhao2024hardware, tuli2023acceltran, you2023vitcod, shen2022salo] have been proposed. Early works, such as A3 [ham20203] and ELSA [ham2021elsa] accelerate computation via approximation techniques. Recent efforts have shifted to jointly optimizing computation and memory. Energon [zhou2022energon] and SOFA [wang2024sofa] adopt fine-grained filtering and FlashAttention-optimized tiling techniques to alleviate memory overhead. However, they still rely on extra sparsity predictors, which become de facto latency and power bottlenecks after sparsification. PADE is the first to explicitly identify and address this issue. It removes the need for external predictors by fine-grained, bit-serial computation, while leveraging efficient bit-level early termination and reuse.
Neural network (NN) accelerator with sparsity. Numerous accelerators [wang2021efficient, hojabr2021spaghetti, wang2018low, gondimalla2019sparten, asgari2020alrescha, deng2021gospa, qin2020sigma, gudaparthi2022candles, fang2022algorithm, hanson2022cascading, lew2022anticipating, li2022ristretto, li2021escalate, li2022accelerating, liu2022s2ta, pavon2021via, rucker2021capstan, sadi2019efficient, walia2021fast, yang2020procrustes, Gon2023Eureka, wu2022sparseloop, wu2023highlight, shin2022griffin] exploit sparsity to accelerate NN inference. General-purpose sparse tensor accelerators [srivastava2020matraptor, mahmoud2020tensordash, kwon2019tensordimm, hegde2019extensor, kanellopoulos2019smash, chen2020tpspmv] support operations on sparse fully connected layers, but most of them target pre-trained, statically sparse weights. By contrast, PADE targets attention dynamic sparsity, which requires on-the-fly prediction, making zero-based sparsity methods ineffective. While some recent works explore activation sparsity [jang2021sparsity] or combine weight and activation sparsity [wu2023highlight, huang2023rm, wang2021dual], they still rely on zero-based sparsity, failing to address the argmax sparsity, which is the PADE targets.
Bit-serial computing accelerators. Prior works [wang2025mcbp, im2023sibia, han2024bitnn, yang2021fusekna, judd2016stripes, albericio2017bit, delmas2019bit, lo2023bit, sharify2019laconic, lee2018unpu, delmas2017dynamic, gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles, lu2021distilling, kam2025panacea, li2025lut] accelerates NNs by exploiting bit-level sparsity [albericio2017bit, delmas2019bit, lu2021distilling, sharify2019laconic, yang2021fusekna] or dynamically reducing bit-width [gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles]. However, these techniques are difficult to apply directly to sparse attention, as they typically rely on offline weight preprocessing, while sparse attention demands real-time, token-wise prediction. Further, the inherently low utilization of bit-level prediction further limits their effectiveness in attention acceleration. In contrast, PADE adopts a lightweight runtime pruning strategy, BUI-GF, and maximizes hardware efficiency via bidirectional bit sparsity and streamlined out-of-order execution.

### v2

@@ V-PU A Systolic Array with 8 $\times$ 16 INT8 PEs;
Benchmarks: We evaluate PADE on several representative Transformer models across NLP and CV tasks. For NLP tasks, Llama2-7B [113], Llama3-8B [38], Qwen7B [6], Bloom1B7 [65] and OPT1B3 [147], for six tasks.These tasks include language modeling (Wikitext-2 (S=2k) [83], Wikilingua (S=2k) [28], Winogrande (S=0.25k) [99]), language understanding (MMLU, S=0.5k) [47], code generation MBPP (S=1k) [4], long context processing dolly (S=15k) [18]. For CV tasks, we choose ViT-L/16 (S=576) [27] and PVT (S=3k) [130] on ImageNet-1k [24] and JFT [110] classification.
Quantization Accuracy. All pre-trained models are sourced from Pytorch [90] and HuggingFace [134]. INT8 baselines derived via post-training quantization, where only the weights and activations (QKVs) are quantized to INT8, while non-linear operators (e.g., softmax) remain in FP16 precision. As shown in Table. II, the INT8 baseline incurs less than a $1\%$ average accuracy drop from FP16, thus confirming its validity.
Hardware Evaluation: Table III lists the hardware configuration of PADE. We implement the RTL design for PADE and utilize Synopsys DC on TSMC 28nm CMOS technology to estimate the logic area and power. The power, area, and read/write bandwidth of on-chip SRAM buffers are estimated through CACTI [84]. Off-chip HBM modeling involves simulating access patterns and row activation under various data layouts (Fig. 22), capturing HBM’s burst behavior. We derive memory latency from Ramulator [62], and estimate IO power following the methodology in [11, 2, 133]. We extract each stage’s cycles by simulating the RTL with Verilator [106], and use a custom cycle-level simulator to evaluate performance.
GPU comparison We benchmark on an Nvidia H100 using SOTA TensorRT-LLM [85] with FlashAttention3 [100]. To exclude the software overhead, we measure execution time with cudaEvent, isolating GPU execution from CPU interference. The GPU is dedicated during testing, and large batch sizes are used to amortize data transfer costs. Non-computational phases are excluded using nvprof. Power is measured via nvidia-smi; dynamic power is computed as the difference between active and idle states. Each experiment is run 2k times, discarding the top and bottom $15\%$ before averaging.
@@ VI-B Algorithm Performance
Fig. 15 (a)(b) compares the accuracy of three representative software-only sparse attention methods, two predictor-free works (SpAtten [116], DTATrans [142]), and PADE. The Sparsity Level denotes the ratio between the computation cost of sparse execution (prediction + actual computation) and that of dense execution. StreamingLLM [138] adopts a static sparsity pattern—retaining only the initial and recent tokens, while MInference [56], and DoubleSparsity [141] rely on runtime sparsity prediction. Four key observations are made: 1) StreamingLLM performs the worst, due to its lack of adaptivity in capturing input-dependent sparsity patterns. 2) MInference improves accuracy by combining dynamic prediction with predefined sparsity patterns but remains limited by restricted pattern diversity. 3) DoubleSparsity introduces a more flexible dynamic sparsity mechanism and integrates channel sparsity to reduce prediction overhead. However, as its prediction computation and memory access cannot be reused in subsequent steps, it suffers from inefficiency, yielding slightly lower accuracy than PADE at the same sparsity level. 4) For SpAtten and DTATrans, comparable accuracy to PADE can be achieved only after fine-tuning. This is mainly because both methods eliminate the predictor by using the previous layer’s attention distribution to guide pruning in the current layer, which introduces significant errors without fine-tuning. In contrast, PADE consistently achieves the best performance across all settings, benefiting from its fine-grained sparsity prediction and efficient reuse of both computation and memory access.
@@ VI-E Comparison to SOTA Accelerators
Workload Balance. Fig. 23 (a) presents a detailed breakdown of execution time across varying numbers of PE lanes to illustrate load balance effects. To better show PADE’s advantages, we compare it against a SOTA bit accelerator, BitWave [104], which leverages bit-flipping to enhance bit-plane sparsity. Since each PE lane integrates multiple bit-serial multipliers, intra-PE stall arises when certain multipliers handle more effective bits. Inter-PE install, in contrast, results from variations in bit sparsity across different key vectors. As the number of PE lanes increase, BitWave suffers from greater intra- and inter-PE imbalance, as it only exploits bit-0 sparsity, which exhibits large variability. In comparison, PADE adopts a more balanced bit sparsity distribution, thereby achieving around $30\%$ higher PE utilization.
@@ VI-F Discussion for Deployment and Scalability
Extension for MXINT. The micro-scale format performs fine-grained quantization along the channel dimension by grouping data into 32-element segments [96]. PADE ensures compatibility by applying group-wise scaling to the bit uncertain interval (BUI) (§IV-A). As shown in Fig. 25 (a), when processing 64-length Q and K vectors, the micro-format partitions them into two 32-element groups and quantizes each group using calibration-derived factors ($\Delta$), resulting in group-wise BUI scaling. As depicted in Fig. 25 (b), PADE achieves compatibility via two steps. ❶ First, it performs bitwise serial multiplication within each 32-length group, with the BUI derived from the strategy in §IV-A. The resulting BUI is then expanded using the quantization factor, eg., $\Delta_{\rm Q1}\Delta_{\rm K1}/\Delta_{\rm A}$. Finally, the max and min BUI values across all groups are aggregated to compute the overall BUI. This method can be extended to dot products of arbitrary length.
Extension for FP formats. Prior studies [103, 137, 149] have demonstrated that the K and V tensors are highly amenable to quantization because the softmax normalization in self-attention naturally suppresses quantization noise. This property enables safe quantization to INT8 or even INT4 with negligible accuracy degradation. Motivated by this, when queries operate in FP format, PADE converts the INT-FP computation into a bit-serial form through exponent alignment, following methodologies adopted in prior works [14, 53, 31]. The resulting bit-serial execution is fully compatible with PADE’s existing processing mechanism.
@@ VI-G Limitations and Future Direction
While PADE significantly advances sparse-attention acceleration by eliminating the sparsity predictor through unified bit-serial stage fusion, several limitations remain for future works: (1) As the sequence length of large language models (LLMs) continues to scale, the need for distributed attention becomes inevitable. Extending PADE to distributed scenarios, especially in emerging wafer-scale architectures [124, 126, 128, 45, 50, 111], remains an open problem. (2) Beyond single-bit granularity, exploring multi-bit stage fusion may provide a more favorable efficiency–accuracy trade-off, which represents a promising direction for future PADE enhancements.
@@ VII Related Works
Efficient Attention Accelerators. Numerous attention accelerators [121, 122, 123, 69, 41, 42, 95, 30, 142, 49, 77, 119, 131, 32, 144, 74, 81, 116, 150, 93, 102, 29, 8, 148, 114, 145, 102] have been proposed. Early works, such as A3 [41] and ELSA [42] accelerate computation via approximation techniques. Recent efforts have shifted to jointly optimizing computation and memory. Energon [150] and SOFA [119] adopt fine-grained filtering and FlashAttention-optimized tiling techniques to alleviate memory overhead. However, they still rely on extra sparsity predictors, which become de facto latency and power bottlenecks after sparsification. PADE is the first to explicitly identify and address this issue. It removes the need for external predictors by fine-grained, bit-serial computation, while leveraging efficient bit-level early termination and reuse.
Neural network (NN) accelerator with sparsity. Numerous accelerators [127, 48, 129, 36, 3, 23, 92, 39, 32, 44, 68, 70, 73, 74, 78, 91, 97, 98, 115, 139, 37, 136, 135, 105] exploit sparsity to accelerate NN inference. General-purpose sparse tensor accelerators [109, 82, 64, 46, 60, 13] support operations on sparse fully connected layers, but most of them target pre-trained, statically sparse weights. By contrast, PADE targets attention dynamic sparsity, which requires on-the-fly prediction, making zero-based sparsity methods ineffective. While some recent works explore activation sparsity [54] or combine weight and activation sparsity [135, 51, 132], they still rely on zero-based sparsity, failing to address the argmax sparsity, which is the PADE targets.
Bit-serial computing accelerators. Prior works [125, 52, 43, 140, 57, 1, 22, 79, 101, 66, 21, 36, 70, 87, 143, 80, 59, 71] accelerates NNs by exploiting bit-level sparsity [1, 22, 80, 101, 140] or dynamically reducing bit-width [36, 70, 87, 143]. However, these techniques are difficult to apply directly to sparse attention, as they typically rely on offline weight preprocessing, while sparse attention demands real-time, token-wise prediction. Further, the inherently low utilization of bit-level prediction further limits their effectiveness in attention acceleration. In contrast, PADE adopts a lightweight runtime pruning strategy, BUI-GF, and maximizes hardware efficiency via bidirectional bit sparsity and streamlined out-of-order execution.